[Torrey, Ian]
General progress notes:
-The three photo diodes required for OFC1 activities are powered.
-All cables to and from the moku/patch panel are set up for OFC1 activities, including PDs, EOM, and DC modulation for the laser.
-1550 path is realigned and can be locked with the laser.
Problem:
The ~900 Hz signal that arrises from having a cable plugged into the DC modulation port of the laser needs to be solved. This is a known problem from when we were in B102 but it seems much worse over here. Here is what I know.
1) Light is blocked. DC modulation port plugged in. ~890 Hz signal observed on PD. This is not optical and must be electrical.
2) Light blocked. DC modulation port is unplugged from the patch panel. Oscillations are gone. To clarify, I am unplugging it here: PXL_20241202_193738276~2.jpg
3) This interferes with the error signal enough to not give a good laser lock.
4) The oscillations can be seen when the channel is DC coupled and light is hitting the PD.
5) Both the laser and the photo detector are plugged into a blue outlet, but not the same blue outlet.
If anyone has any insights on this let me know. I am curious to see if the PD is plugged into a white outlight if this makes a difference. Lee has suggested a type of circuit to fix this potential grounding issue in the past although I am blanking on the name of it.
After some frustrating moku glitches and then some realignment, the second output filter cavity now has 775 light. Pending both 775 paths, and 1550 in OFC2, we should be ready for filter cavity science again.
[Sander, Daniel]
We routed a USB Extension Cable from the vacuum oven in B110 through B111D to the B111B control room. The cable trays above the mobile clean rooms in B111B were very hard to access. Maybe we should get some sort of cheap grabber tool from Amazon.
The cables make it to the computers in B111B but just barely. A ~6 ft USB extension cable would ensure there is plenty of slack in the system.
The computer recognized that something was plugged in, so it seems the extension cables work. Next week, I will finish setting up the vacuum oven software that I downloaded. I previously got this software to work on my laptop.
The computer is able to communicate with the vacuum oven after very minimal software setup.
I added another 50" USB extension so that there isn't tension on the cables. This is ~45" too long, but the extra cable is not in the way and the oven is able to communicate with the computer and vice versa.
I have checked all visible lab ethernet ports and compiled the following csv file that shows all ports are now on the same subnet. All ipv4 addresses pulled by my laptop returned 192 as the subnet address.
Thus we should be able to use dns on all instruments and connected devices now.
There may need to be changes in the active DNS naming on Pihole to correct any instruments that were previously on a switch that was not on the correct subnet.
Please take a look at the attatched csv and you can access my code for grabbing ipv4 addresses and storing them in a csv as such here.
(To run code open a terminal, navigate to file, enter "python3 ipv4_grabberV2.py" and follow the instructions. You may need to install Pandas if not already installed)
In my sporadic work-from-home I'm fixing up some of the computing infrastructure.
We now have incremental backups for the files in the nextcloud. If you lose something particularly important, we can recover them. I haven't made a means to do this as a user yet (namely, read-only access), but might at some point.
It will not back up files larger than 1Gigabyte - beware that detail.
Also, if you have long-lived services that you want to monitor (like backups). You can use the https://healthchecks.mccullerlab.com/ service. It is watching all of the backups of the web services and nextcloud. It is running the same service as at healthchecks.io. It just watches for pings to http address on a schedule.
This will be used to back up and monitor the various lab machines as well. That is nearly set up.
In order to connect the Power Recycling Vacuum Cube to the Central Vessel, the heights of the center flanges must be close (more tolerance with the use of bellows). I measured the distance between the bottom of the 10" side length vacuum cube and the custom aluminum base to be between 0.97" and 0.99", with an average of 0.976" (nominal thickness is 0.97"). I measured the custom aluminum base to be between 0.5" and 0.52" (nominal thickness of 0.50"). We will likely need a 0.01"-0.02" sheet to bring up the cube bottom to 1.5" above the optics table. We have shim stock to do this.
[Ian, Jeff, Sander, Daniel]
Ian and I removed the 10" to 4.5" zero length reducer from the vacuum cube that was the "elevated" Holometer bend cube so that we could loosen a screw so that we could later get the "elevation part" off that elevated the cube. We then installed a new copper gasket and tightened the zero length reducer "normally" to 34 Nm with 1.75" long screws (I've started testing before removing the old base to make sure these screws go in all the way.
Ian, Sander, and I then lifted the cube and loosened the last screw holding the elevation part to the base flange and removed this part. We then set the cube down and tilted it over.
Jeff and I replaced the base flange with the custom flange I made and custom aluminum base to secure the cube to the table. I tightened all the 1.75" long screws to 34 Nm.
This cube can be flipped over and placed on the table.
I cleaned stainless steel PH0 phillips, PH2 phillips, and flathead screwdrivers for use with components that will go in vacuum. I also cleaned a stainless steel pair of scissors. The cleaning process was the standard stainless steel cleaning procedure.
[Jeff, Ian, Sander, Daniel]
We tipped over a 10" vacuum cube like we did before so that we could replace the bottom CF Flange with one that I tapped.
We vented the vacuum oven from ~0.2 atm to 1 atm and grabbed one of the custom flanges that I tapped. We attached the CF flange to the cube with a new method courtesy of Maty Lesovsky: screw in a few screws at the bottom of the flange, tip the flange so the bottom is touching the cube and the top is exposed, then drop in the copper gasket. We took off the screws and rotated the flange to ensure the gasket was in the correct place.
We tried to use silver-plated 1.75" long 5/16-24 screws to secure the flange to the cube, but we couldn't thread them in deep enough so that the head (and washer) made contact with the flange. We therefore attached the 1.75" screws with hex heads through the custom base I had made to secure the flange and replaced the silver plated 1.75" long screws with silver plated 1.5" long screws. We could use the 1.75" screws with hex heads becuase the custom part I made puts the heads ~0.25" from the flange. I noticed that I partially stripped some of the silver plated 1.75" screws.
After screws of an appropriate length were inserted, I tightened them in a star pattern with a torque wrench, starting at 13 Nm and ending at 34 Nm, with increments of 6.8 Nm. By the time I went around the flange at 34 Nm a few times and no screw was further tightened, the flange was very close or touching the cube everywhere.
See this post for images of a similar procedure.
[Ian, Sander, Daniel]
This cube has been tipped right side up and is ready to be moved onto the optics table.
I have changed the name of the TP-link subnet access to 'B111net'
Check lab secrets for the password
not "pLANck" :(
[Ian, Lee]
It was suggested that I have a gentler roll off to the FBNS FOM. I removed two polls from the roll off so it is now gentler. This should help the solver. Attached is the plot with the simplified FOM (the one labeled "Down sampled FOM ASD"), the fit from the data2filter, as well as my hand fit which is shown in green. This FOM is normalized and the normalized FOM is what is actually attached to the model. The normalized FOM is shown in the second attachment. Also I have attached the .yml file for the pre-normalized FOM.
Note in the first plot the black dots and red crosses represent the positions of the zeros and polls. just to give you an idea of where they were.
[Ian, Torrey, Sander]
We moved the sleds for the rubidium vapor cell locking experiment as well as the sled that holes the SHG optics for the 1560 nm to 780 nm conversion to the north most small table in the atoms lab (B11A). We disconnected all of the cables in B102 and moved the sleds as well as the 780 nm laser (McCuller Lab SN: S2400003) and the laser current and temp controllers. We also moved the temp controller for the vapor cell. I haven't moved any of the cables for any of the instruments of hooked anything up yet. I plan on trying to do that sometime on Friday. Then I will briefly turn on the laser to make sure it works.
Because we moved the laser we need to update the standard operating procedures (SOP) for the laser so that it refers to the new lab. We also need to update the Caltech property tag location. The Caltech property tag is (C0E000083455). I have already updated the wiki for the location. As well as added a link to the SOP there.
[Ian, Sander, Torrey]
As part of the Watson Lectures at Caltech, we gave a tour of the lab spaces and gave a brief talk to some students from South Pasadena High School before an Athenaeum dinner with them and attending the lecture by Prof. Chatziioannou. We showed them a video describing gravitational waves before we gave a brief talk on exactly how GQuEST is different from LIGO and what Sander and I both work on. We then gave them a tour of the lab and showed them inside of one of the laser cavities. They were very interested in lasers so we also discussed some of the laser work that is going on in the LIGOX group and LIGO in general. It was a good event.
We should invest in a model interferometer that we can use to demonstrate how it works to students and the public. It would have been helpful.
Continuing realignment of sleds in B111B.
-Powered SHG for the first time. Note the SHG is able to be controlled via USB on the computer, and I have a USB hub near enough for this to reach. It is not plugged in currently while I decide on a final spot for the hub. Turned on and brought to temperature via the buttons on the temperature controller. Another note that we may want to double check the optimal temperature down to the 0.001 of a degree.
-Frequency double light is coming out of the SHG. Touched up the input and output polarizations for the SHG.
-Want to realign each AOM path + fiber. Starting with AOM fiber path 4 (as labeled on the optical fiber). Powered the AOM with the patch panel + homemade chassis successfully. Replaced the mirror after the quarter waveplate to a curved f = 50 mm, ROC = 100mm mirror. The once shifted light is sent back through the aom and then fiddle with the alignment until the twice shifted beam pops up. Efficiency is 208 uW input to the AOM, 58 uW of twice shifted headed to the fiber. This is on par with previous alignments.
-Realigned into the fiber. Output is 45 uW (78% efficiency). OFC 1 has both wavelengths of light now and is ready to be realigned.
I did some organization of parts for in vacuum and working around vacuum. Into the short lista cabinet in the north west corner of B111B, I put screws to attach CF flanges, anti-static bags to hold open vacuum parts (that are already wrapped in foil), CF Blank Flanges, CF reducer flanges, and other vacuum parts with CF Flanges.
Into a white mobile shelving unit, I put mirror mounts, optics posts, and clamps. I also put silver plated screws for in vacuum use into some stainless steel bento boxes that I cleaned. One box has silver-plated 1/4-20 hardware with socket heads and another box as silver-plated #8-32 hardware and 1/4-20 button head screws. I wrapped each box in UHV foil and put each in its own anti-static bag and labeled the bag with its contents. Please see the attached photo for the exact hardware type. The labels and the empty bags in the photo are placed how the bento boxes are organized.
Some notes on preferred networking setup now that I see machines going up with posts:
https://mccullerlab.com/logs/lab/index.php?callRep=11985
and https://mccullerlab.com/logs/lab/index.php?callRep=11998
The lab LAN uses a specially configured subnet and DHCP. You should not need static IP addresses, since the DCHP server remembers and maintains the assigned addresses.
The subnet settings are:
192.168.248.1/21 (equivalent \21 mask is 255.255.248.0)
The gateway is 192.168.248.1
broadcast address is: 192.168.255.255 (can be calculated from the above with the tool https://jodies.de/ipcalc?host=192.168.248.1&mask1=21&mask2=21)
the following addresses are supported:
HostMin: 192.168.248.1 HostMax: 192.168.255.254 Hosts/Net: 2046
The synology is running the internal DNS server a internal LAN address 192.168.248.15 so it should be included in your resolve.conf or equivalent. DHCP currently assigns it as the secondary DNS server. The primary is a preferred CIT server. We could make the internal DNS the primary, but may have bigger issues if/when it goes down.
if you assign static IP's, please only assign them to addresses already registered by the DHCP to the particular MAC address of the interface.
The LIGO outreach interferometer got upgraded to a real breadboard and ThorLabs parts. They needed a way to mount their laser pointer. I used the plastic mount from the old mount as well as some of the half-inch hardware to mount it.
This equipment includes:
We (meaning LIGO) should really get a Thorlabs laser that is brighter and more reliable than a laser pointer. One suggestion is this the Thorlabs CPS532-C2 with the DS5 power supply so that it can be run off of a USB A battery pack. The Thorlabs MK11F would be perfect for mounting it on a breadboard. Also a viewing screen like the Thorlabs EDU-VS2 would be perfect. Thorlabs also offers a Michelson interferometer educational kit that has lots of the components that would be needed but it includes too many extras and the breadboard is overkill for this application.
[Lee, Torrey, Daniel]
On Friday, we adjusted the loop shaping on the filter cavity to try and make the lock better. With a 1/f loop, it will lock, but goes between flickering out of lock, to oscillating as the gain is increased.
The new shaping uses the "slow" output of the "laser lock box." It shapes the fast output to have a 10Hz - 200Hz boost (pole - zero), giving 20x gain at low frequencies. The overall 1/f is then put at 1Hz in the slow path. We are able to get a UGF of 1kHz, with 30-40deg remaining. We lose 90deg by 2kHz from either the actuator or the Moku, so 1kHz is around the max UGF we can currently get.
Even with that loop, we get dropouts while talking loudly or while the fans are on. Problematic. That mean the loop can supress the peak length noise below 1.5um / 2 / (2 * finesse). Our current input coupler is 1%, so our finesse is around 2pi / 0.01 = 600 (it might be as low as 300, depending on the losses in the cavity). So our peaks/RMS are around 0.5nm - 1nm of noise.
+1 to the piezo transfer teams. both 11380 and 11373 have the transfer function data along with the plots. It looks like all of the phase we are losing in the cavity lock is from the piezo (though could still be from Moku). We'll need to fit and at least slightly invert this actuator transfer function to make a loop above 1kHz UGF. The fit can help tell us how much of the phase is from the piezo roll-off.
We are going to need it - the current loop & cavity has marginally too much noise, with an RMS around the linewidth. That is going to be a problem once we increase the finesse of the cavity. So we'll need to reduce the noise and/or improve the loop a respectable amount.
our voices are driving 100-300Hz, so we need gain there. If we can get the UGF up as high as 3kHz, then we should also be able to move the boosts up to 600Hz or so. That should give 9x more loop gain and cover much of the need.
options:
frequency lock - good for acquiring and transitioning to aggressive lock, but not an option for operating the experiment
optimal control - we can determine how good we can do with what we have
fitting/inverting actuator - probably needed to get UGF higher. May vary with time in bad ways
sound attenuation or active noise cancellation - may be possible and should look into engineering this.
We should determine if the sound driving is from pushing the mirrors or from the index fluctuation as the pressure modulates. 11290 has calculations about the pressure pushing on a 1kg mass. My worry is that the small alignment adjust set-screws in the MKS flexture mounts are fairly compliant. Could sound drive them 1nm?
My guess is that the light going through 2m of air is the culprit.
The sound pressure of 40dB should be around 2000uP. Air pressure is 101kPa. The index of air is 1.000293
2m / (1.000293**2) * 2000e-6Pa / 101kPa * (1.000293 - 1) = 1.2e-11m
So that is a little small, but if our sound is at 80db, which would be quite loud, then it would be enough to cause this.
For the mirror motion, the mirror mounts are about 1.5in square. On the pressure level that is
(.0254m * 1.5)**2 * 2000e-6Pa = 3e-6N
of the mirror+panel is about 35g (very rough estimate). For a free mass motion that is
1 / ((2 * np.pi * 100Hz)**2 * 0.03534291735288516kg) * 3e-6N = 2e-10m which is getting in the ballpark with 60db of noise rather than 40db. The mirrors aren't free though.
and now I'm curious how that is against the spring constant of the tiny screw.
Supplement to Lee's post:
1) IMG_0005.png is the scheme used to achieve lock in multi-instrument mode.
2) IMG_0004.png is the closed loop gain while the cavity is locked. Note the UGF of ~1kHz as Lee mentioned.
3) IMG_0009.png is the slow controller loop.
4) IMG_0008.pngis a spectrum for f<10kHz. Note the resonance at ~3kHz. This may be the culprit and will be investigated.
5) IMG_0007.png is the fast controller loop just being used as a low frequency boost.
6) IMG_0010(1).png is the overall structure of the laser lock box as a reference for people that build this system in the future.
Quick test to see if the air theory is likely. IMG_0011.png is cavity locked, fans off, no noise emitting as a control. IMG_0014.png is cavity locked, emitting a 1kHz sound from my phone about 1.5 meters away from the cavity with the lid off (reference trace) and lid on (active measurement). Noise is definitely muted. Air theory seems plausible.
Flexure Mount spring constant: E*A/L = 200 GPA * pi * 0.035 in ^2 / 0.1 in = 2*10^8 N/m
Resonant frequency: 75 kHz
RMS displacement: 2 * 10^-14 m for 40 dB of sound
The small piece of metal that flexes on this mount is 0.020" thick and 1.5" long according to SolidWorks (this seems about right experimentally). This is a cross section 10x of the screw, so this also doesn't explain the acoustic pick-up.